# Nonlinear integral extension of PID control with improved convergence of perturbed second-order dynamic systems

### v1

@@ III-C Extension by an actuator dynamics
Then, the extended plant (23), (22) is controlled in the same way by (8), provided the resulted Lur’e system equally fulfills the necessary assumptions i)-iv), cf. with section II-B.
For the extended plant (23), (22), the corresponding input-output loop transfer function results then in
@@
Since neither the structure of a feedback system nor the nonlinearity (11) change by (22), the same circle criterion as given in section II-B is applicable to (24). Respectively, the parametric inequality (12) of Theorem 2 should modify, so that to take into account also the polynomial coefficients of $F(j\omega)$. Since both the absolute stability and disturbance rejection for (24) follow the same lines of argumentation as in sections III-A, III-B, and a particular actuator dynamics $F(s)$ is rather application specific, we omit a detailed development here. At the same time, it is worst emphasizing that the experimental case study provided below in section IV has the second-order system plant additionally affected by such first-order actuator dynamics $v(s)=\kappa(s\mu+1)^{-1}u(s)$, with a gain factor $\kappa$ and (not negligible) time constant $\mu>0$.
IV EXPERIMENTAL EVALUATION
The proposed extended nonlinear PID control given by (8) is experimentally evaluated and additionally benchmarked with a standard PID control (2). In addition, the corresponding PD control (i.e. $c=0$) is also evaluated for highlighting the level of perturbations in the system at hand and, this way, disclosing necessity of the integral control action. The common control parameters are assumed the same, while all parameter values are summarized in Table I.
TABLE I: Control parameters
control \  params. $a$ $b$ $c$ $d$ $e$
@@ IV EXPERIMENTAL EVALUATION
The experimental laboratory setup used for evaluation of all three controllers is depicted in Fig. 7.
@@
The motion system consists of a voice-coil driven actuator with 1DOF and contactless measured output position $y(t)$. The latter contains a relatively hight level of the sensing noise. The robust finite-time second-order sliding-mode differentiator [25, 26] is used for obtaining $\dot{y}(t)$. The sampling frequency of the implemented real-time control is set to 10 kHz. A more detailed description of the experimental system, including the identified system parameters, can be found in [18].
Further it is worth noting that: i) an additional first-order actuator dynamics (due to electro-magnetic behavior of the voice-coil motor) is in place, and ii) the unknown perturbation $\sigma(t)$ is mainly due to the gravity, nonlinear friction, and force ripples of the anchor-stator interaction.
The measured output response to the step reference $r=0.008$ m, that is applied at time $t=1$ sec, is shown in Fig. 8 for the nl-PID, PID, and PD controllers.
@@
The absolute output error $|\epsilon|\equiv|r-y|$ of the nl-PID and PID controllers is additionally visualized on the logarithmic scale in Fig. 9. One can recognize that for $|\epsilon|<0.0001$ m, both controllers have to deal with a sensing noise so that the residual control errors are similar. At the same time, the convergence performance is well distinguishable for the time range between 1.18 and 1.7 sec. The transient behavior (including over- and undershoots) at the time between 1 and 1.18 sec is well comparable.
@@
V CONCLUSIONS
In this paper, a novel nonlinear extension of the PID feedback control (denoted by nl-PID) is introduced. The nonlinear extension is proposed for integral control action that allows improving convergence performance in presence of the matched unknown perturbations in the second-order systems. The nonlinear integral control part does not change the system structure and is continuous in time and Lipschitz in the system output variable, so that the absolute stability by means of the circle criterion turned out applicable and sufficient for analysis. The nl-PID control has five design parameters in total, while keeping the same meaning and tuning rules for the proportional, integral, and derivative feedback gains. The global asymptotic stability of the nl-PID control was shown for the case of constant perturbations. For the case of Lipschitz perturbations, an ultimately bounded output error can be guaranteed by consideration in frequency domain. Beyond the analysis and numerical examples, an experimental case study is demonstrated for benchmarking the standard PD, PID, and proposed nl-PID controls.

### v3

@@ 3.4 Extension by an actuator dynamics
Then, the extended plant (23), (22) is controlled in the same way by (8), provided the resulted Lur’e system fulfills equally the necessary assumptions i)-iv), cf. with section 2.2.
For the extended plant (23), (22), the corresponding input-to-output loop transfer function results in
@@
Since neither the structure of a feedback system nor the nonlinearity (11) change through (22), the same circle criterion as given in section 2.2 is applicable to (24). Respectively, the parametric inequality (12) of the Theorem 2 should modify, so that to take into account also the polynomial coefficients of $F(j\omega)$. Since both the absolute stability and disturbance rejection for (24) follow the same lines of argumentation as in sections 3.1, 3.2, and a particular actuator dynamics $F(s)$ is rather application-specific, we omit a further detailed development here. Also note that the controller design procedure, summarized in form of the step by step recommendations given above in section 3.3, is correspondingly extendable by the actuator dynamics $F(s)$. Here we recall that the actuator dynamics $F(s)$ will generally reduce the characteristic stability margins of the closed-loop (cf. e.g. 3, 4 for details), and thus require a PID control design to be more conservative regarding the feedback gain values.
At the same time, it is worst emphasizing that the experimental case study provided below in section 4 has the second-order system plant affected additionally by such first-order actuator dynamics $v(s)=\kappa(s\mu+1)^{-1}u(s)$, with an actuator gain factor $\kappa$ and a (not negligible) time constant $\mu>0$.
4 EXPERIMENTAL EVALUATION
The proposed extended nonlinear PID control given by (8) is experimentally evaluated and additionally benchmarked with a standard PID control (2). In addition, the corresponding PD control (i.e. $c=0$) is also evaluated for highlighting the level of perturbations in the system at hand and, this way, disclosing necessity of the integral control action. The common control parameters are assigned to be the same, while all parameter values are summarized in Table 1. Note that all control gains are normalized by the overall gain $\kappa$ of the double integrator plant, cf. (14), so that $a^{*}=a/\kappa$, $b^{*}=b/\kappa$, $c^{*}=c/\kappa$, $d^{*}=d/\kappa$.
Table 1: Control parameters
control \  params. $a^{*}$ $b^{*}$ $c^{*}$ $d^{*}$ $e$
@@ 4 EXPERIMENTAL EVALUATION
The experimental laboratory setup used for evaluation of all three controllers is depicted in Figure 7.
@@
The motion system consists of a voice-coil driven actuator with one translational degree of freedom, which has the contactless measured output position $y(t)$. The latter contains a relatively hight level of the sensing noise. The robust finite-time second-order sliding-mode differentiator 17, 20 is used for obtaining $\dot{y}(t)$. The sampling frequency of the implemented real-time control is set to 10 kHz. A more detailed description of the experimental system, including the identified system parameters, can be found in 23.
Further it is worth noting that: i) an additional first-order actuator dynamics (due to electro-magnetic behavior of the voice-coil motor) is in place and ii) the unknown perturbation $\sigma(t)$ is mainly a combination of the gravity, nonlinear friction, and force ripples of the anchor-stator interaction.
The measured output response to the step reference $r=0.008$ m, that is applied at the time $t=1$ sec, is shown in Figure 8 for the nl-PID, PID, and PD controllers.
@@
The absolute output error $|\epsilon|\equiv|r-y|$ of the nl-PID and PID controllers are additionally visualized on the logarithmic scale in Figure 9. One can recognize that for $|\epsilon|<0.0001$ m, both controllers have to deal with the sensing noise so that the residual control errors are similar. At the same time, the convergence performance is well distinguishable for the time interval between 1.18 and 1.7 sec. The transient behavior (including over- and undershoots) at the time between 1 and 1.18 sec is well comparable.
@@
5 CONCLUSIONS
In this paper, a novel nonlinear extension of the PID feedback control (denoted by nl-PID) is introduced. The nonlinear extension is proposed for integral control action that allows improving convergence performance in presence of the matched unknown perturbations in the second-order systems. The nonlinear integral control part does not change the system structure and is continuous in time and Lipschitz in the system output variable, so that the absolute stability by means of the circle criterion turned out applicable and sufficient for the analysis. The nl-PID control has five design parameters in total, while keeping the same meaning and tuning rules for three of them – proportional, integral, and derivative feedback gains. The global asymptotic stability of the nl-PID control was shown for the case of constant perturbations. For the case of Lipschitz perturbations, an ultimately bounded output error can be guaranteed by consideration in frequency domain. Beyond the developed analysis and numerical illustrative examples, an elaborated experimental case study is demonstrated for benchmarking the standard PD, PID, and proposed nl-PID controllers.